# Beyond Aesthetics: Cultural Competence in Text-to-Image Models

### v5

@@ 5 Evaluating Cultural Diversity
We seek to assess the cultural diversity of T2I outputs across different seeds as a way to measure the model’s intrinsic latent space cultural diversity (Xu et al., 2024). For instance, a model capable of generating a diverse array of cultural artifacts across a range of seeds demonstrates the cultural richness of its learned representations. A more detailed note on our motivation for seed variation is outlined in Appendix C. For this, we focus on under-specified prompts (Hutchinson et al., 2022) - prompts that elicit the generation of diverse cultural artifacts (e.g. "Image of tourist landmarks") rather than specific objects (e.g. "Image of Eiffel Tower"). We then ask: What is the global inclination and geo-cultural diversity of the generated cultural artifacts for prompts that mention just a concept? The CUBE collection encompasses a large set of culturally salient artifacts, which serves as a broad-coverage grounding for diversity evaluation.
@@ 5.1 Cultural Diversity (CD)
Existing works that focus on visual diversity use LPIPS (Zameshina et al., 2023) and Coverage (Hall et al., 2024), based on image embeddings. However, these metrics are not directly applicable in our case, as the similarity here may be attributed to color, texture, spatial orientation, and other visual aspects of the images. Measuring the cultural diversity of text-to-image (T2I) models requires a nuanced approach that accounts for both the variety of generated cultural artifacts and the quality of images generated from text prompts. To address this, we introduce Cultural Diversity (CD), a new T2I evaluation component leveraging the quality-weighted Vendi Score (Nguyen & Dieng, 2024).
@@ 5.1.2 Incorporating Quality: Quality-Weighted Vendi Scores
While Vendi scores excel in measuring diversity, they treat all items equally without considering individual quality. In the context of T2I outputs, however, it is crucial to account for the quality of the generated images conditional on text prompts. We therefore rely on quality-weighted Vendi scores (qVS) (Nguyen & Dieng, 2024) that extends VS to account for the quality of items in a given collection. qVS is defined as the product of the average quality of the items in the collection and their diversity,
@@
We employ the HPS-v2 metric (Wu et al., 2023) as the $s(\cdot)$ function to score the quality of T2I outputs. HPS-v2 provides a quality score $s\in[0,1]$ for of an image conditioned on text prompt. HPS-v2 is trained on 790k human preferences across key aspects such as faithfulness, aesthetics, and realism of generated images, making it suitable proxy to measure image quality in our case. We leave exploration of other quality metrics that can also accommodate salience of generated artifacts for future work.
@@ 5.1.3 Desirable Properties of $q\overline{\mathrm{VS}}$
We state above three core properties of $q\overline{\mathrm{VS}}$ that makes it suitable for measuring cultural diversity in T2I models: 1) prioritizes collections with higher-quality items when other factors are equal, 2) penalizes the duplication of elements, and 3) exhibits flexibility in capturing various aspects of diversity through the selection of an appropriate similarity kernel. The proof of the quality-awareness property is immediate following the definition of $q\overline{\mathrm{VS}}$. See Appendix E for a proof of duplication scaling. While the kernel generalizability is intuitive, we demonstrate its utility in measuring different aspects of cultural diversity in Section 5.2.2.
@@ 5.2 Experimental Setup
We seek to analyze the global geo-cultural diversity of generated artifacts for under-specified prompts (Hutchinson et al., 2022) that simply mention the concept, such as "Image of traditional clothing.". These prompts serve as great test-beds to analyse model’s intrinsic cultural diversity. We briefly describe the setup below and elaborate on finer details in Appendix H.3. We further analyze within-culture diversity (What is the diversity of cultural artifacts produced by model for different cultures?) for Imagen 2 in Appendix D.
5.2.1 Prompting and Seeding
To evaluate the geo-cultural diversity of text-to-image models, we employ a straightforward prompting strategy where the prompt consists solely of the target concept (e.g., "Image of landmarks"). This allows us to measure both the model’s intrinsic cultural inclination and the breadth of its generated artifacts. Table 2 (right half) presents some examples of these prompts to measure geo-cultural diversity. To minimize the variance due to prompt wording and seed selection on our results, we utilize five prompt template variants and ten different seed batches for each prompt. Each seed batch consists of images generated from eight consecutive seeds. We limit our diversity analysis to eight images, aligning with the typical output of image-generation APIs that present users with four or eight images per prompt. This results in the analysis of 400 images per prompt (5 prompts x 10 seed batches x 8 images/batch). Hence, our reported scores represent the mean over 50 repetitions. For more details on our prompting and seeding strategy, please refer to Appendix H.1.
5.2.2 Computing Cultural Diversity
Mapping generated images to cultural artifacts.
To compute the cultural diversity of the generated images, we first map each image to its most closely resembling artifact from CUBE-CSpace. This mapping process is crucial for grounding our analysis in real-world cultural artifacts [^3] We acknowledge that not all text-to-image generated images perfectly represent real-world cultural entities.. Given that our prompts simply mention global concepts, this mapping allows us to associate each generated image $I$ with its corresponding continent $c$, country $r$, and artifact name $a$, i.e., $I\in\{c,r,a\}$. We develop an automated method leveraging GPT-4-Turbo and mSigLIP (Zhai et al., 2023) based retrieval techniques to perform this mapping and discuss potential limitations, as well as additional details, in Appendix H.2.
@@ 5.2.2 Computing Cultural Diversity
  • Artifact-level diversity: $w_{1}=0$, $w_{2}=0$, $w_{3}=1$. This ignores geographical associations and measures diversity purely based on distinct artifacts.
@@
  • Uniformly weighted diversity: $w_{1}=1/3$, $w_{2}=1/3$, $w_{3}=1/3$. This provides equal weight to all three forms of similarity.
Further details regarding the computation of cultural diversity scores as well as justification for the usage of HPS-v2 as the quality score can be found in Appendix H.3.
5.2.3 Models Evaluated
@@ 5.3 Results
Our analysis reveals that when prompted with under-specified prompts mentioning only general concepts (Table 2), current T2I models tend to generate artifacts that lack comprehensive geographical representation. This finding aligns with previous observations (Basu et al., 2023), suggesting a bias towards well-represented and popular countries. Figure 8 illustrates this tendency for the concepts of cuisine and landmark, respectively.
@@
Table 5 presents the results for both the average quality score (q) and the diversity component ($q\overline{\mathrm{VS}}$) across different kernels. Playground and Imagen generally achieve the highest quality scores based on the HPS-v2 metric. As anticipated, models exhibit the lowest diversity for $w_{1}=1$, $w_{2}=0$, $w_{3}=0$, which considers only continent-level similarity, due to the limited number of continents. Conversely, $w_{1}=1$, $w_{2}=0$, $w_{3}=0$, focusing solely on artifact diversity, yields the highest scores, reflecting the wide array of potential cultural artifacts. In terms of overall performance, Imagen 2 consistently demonstrates the best $q\overline{\mathrm{VS}}$ scores across different kernels for the Cuisine and Art concepts, whereas SDXL obtains the highest scores for Landmarks concept.
Table 6: CD scores across models and concepts using various similarity kernels averaged over 50 repetitions. Imagen 2 (IM) performs best for Cuisine and Art, while SDXL performs best for Landmarks. Importantly, even the best scores are relatively low, indicating significant room for improvement in the global geo-cultural diversity of T2I outputs.
@@
Table 6 shows the cultural diversity ($q\overline{\mathrm{VS}}$). Importantly, we note that even the highest scores achieved are still relatively low, remaining far from the maximum score of 1. This underscores a crucial observation: while some models demonstrate a degree of geo-cultural diversity, there is substantial room for improvement across the board. Current T2I models are, in essence, falling short of encompassing the true breadth and richness of global geo-cultural diversity.
6 Correlation Analysis
@@
7 Discussion
To the best of our knowledge, CUBE is the first large-scale cultural competence benchmark for text-to-image models. From our investigations so far, one clear finding stands out: there is yet significant headroom for improvement of global cultural competence in the current generation of text-to-image models — both in terms of awareness and diversity. This seems especially true for the Global South, highlighting the urgency of the need for comprehensive and informative cultural competence testing frameworks. Towards that goal, we intend to make the CUBE datasets and metric implementations publicly available on publication of this paper, and encourage its adoption and expansion by the multimodal generative AI community.
Design of this benchmark required many challenging decisions that needed careful thought. The increased coverage of our benchmark as a result of largely automated, scalable approaches - built on curated data sources and existing model capabilities still comes with its own limitations. For instance, existing structured knowledge bases such as WikiData are known to have inherent cultural biases reflecting disparities in global distribution of knowledge production (Callahan & Herring, 2011; Bjork-James, 2021). Therefore, it is important to note that our approach of expanding coverage using existing knowledge bases should be complemented with community based and participatory approaches for richer socio-cultural representation (Alonso Alemany et al., 2023; Dev et al., 2024). Notably, the World Wide Dishes effort (Magomere et al., 2024) builds a dataset of images representing dishes from around the world through a community-led effort that complements our dataset that relies on existing knowledge bases.777www.worldwidedishes.com
Furthermore, even with significant filtering and completion, we expect our data to be noisier than other methods. We have not yet explored the potential utility of our curation method, nor of the dataset itself, to empower other research on evaluating or improving cultural competence in generative models in general. Many aspects of the curation process for this benchmark are automated, due to the large scale of the problem — e.g., we rely on image similarity scores and LLM-based selection flows to ground generated images to a specific cultural artifact. This approach risks ingraining biases that may already exist within these tools into the benchmark construction itself. The process of mapping artifacts to country/continent may also introduce biases since the annotator VLM itself may not be aware of several cultural artifacts around the world. On the other hand, human annotation to measure faithfulness and realism is also quite challenging and subjective, as annotators may not be aware of the multitude of representations of their own culture. Even beyond cultural knowledge, different cultures may also have different standards for realism of their images — which could result in mis-calibrated results obtained from human annotations, making it hard to compare across different cultures (e.g., see Table 13).
8 Conclusion
We introduced CUBE, a new T2I benchmark that assesses the cultural competence of T2I models along two crucial dimensions: cultural awareness and cultural diversity. We presented a scalable methodology with a potential to be scaled beyond eigth countries and three concepts considered in this work. Furthermore, we proposed a novel T2I evaluation component: cultural diversity (CD) and measured it using the quality-aware Vendi score. Our comprehensive human evaluation reveals substantial gaps in cultural awareness across cultures and concepts, as well as the gaps in the geo-cultural diversity of model generations. Our correlation analysis reveals a noteworthy trend: while faithfulness and realism exhibit a moderate positive correlation, suggesting they can be improved in tandem, cultural diversity remains weakly correlated to these metrics. This underscores the need to explicitly prioritize diversity as a key objective during T2I model development. In the future, we aim to expand the resource to a wider array of concepts and countries, and also explore how well this technique extends to other cultural lenses. By highlighting existing limitations in cultural competence of T2I models, we believe our work contributes to a critical dialogue surrounding the development of truly inclusive generative AI systems.
9 Ethical Considerations
10 Contributions
11 Acknowledgements
Appendix A Additional details of CUBE
Below we provide the CUBE-1K dataset breakdown as well as the WikiData nodes selected as root nodes for CUBE extraction
A.1 CUBE-1K statistics
@@ A.2 Wikidata
Appendix B Detailed Human Annotation Questions
@@
Table 11: Inter rater reliability for the 3 annotation tasks described in Section B and for all the rater pools across the different locations of our study. For the Cultural Relevance question, we report the observed majority agreement. For both the Faithfulness and the Realism questions, we report the Krippendorff’s $\alpha$.
@@
Appendix C Background on seed variation in T2I models

### v6

@@ 5 Evaluating Cultural Diversity
We seek to assess the cultural diversity of T2I outputs across different seeds as a way to measure the model’s intrinsic latent space cultural diversity (Xu et al., 2024). For instance, a model capable of generating a diverse array of cultural artifacts across a range of seeds demonstrates the cultural richness of its learned representations. A more detailed note on our motivation for seed variation is outlined in Appendix H. For this, we focus on under-specified prompts (Hutchinson et al., 2022) - prompts that elicit the generation of diverse cultural artifacts (e.g. "Image of tourist landmarks") rather than specific objects (e.g. "Image of Eiffel Tower"). We then seek to answer: What is the geo-cultural diversity of the generated cultural artifacts for prompts that mention just a concept?. We further study the within-culture diversity in Appendix (J) and perform a correlation analysis of cultural diversity with existing metrics (in AppendixB).
@@ 5.1 Cultural Diversity (CD)
Existing works that focus on visual diversity use LPIPS (Zameshina et al., 2023) and Coverage (Hall et al., 2024), based on image embeddings. However, these metrics are not directly applicable in our case, as the similarity here may be attributed to color, texture, spatial orientation, and other visual aspects of the images. Measuring the cultural diversity of text-to-image (T2I) models requires a approach that accounts for both the variety of generated cultural artifacts and the quality of images generated from text prompts. To address this, we introduce Cultural Diversity (CD), a new T2I evaluation component leveraging the quality-weighted Vendi Score (Nguyen & Dieng, 2024).
@@ 5.1.2 Incorporating Quality: Quality-Weighted Vendi Scores
While Vendi scores measure diversity, they treat all items equally without considering individual quality. In the context of T2I, however, it is crucial to account for the quality of the generated images conditional on text prompts. We therefore rely on quality-weighted Vendi scores (qVS) (Nguyen & Dieng, 2024) that extends VS to account for the quality of items in a given collection. qVS is defined as the product of the average quality of the items in the collection and their diversity,
@@
We employ the HPS-v2 metric (Wu et al., 2023) as the $s(\cdot)$ function to score the quality of T2I outputs. HPS-v2, trained on 790k human preferences, provides a quality score $s\in[0,1]$ for an image conditioned on text prompt, making it suitable proxy to measure image quality in our case. We leave exploration of other quality measures of salience of generated artifacts, for future work.
@@ 5.1.3 Desirable Properties of $q\overline{\mathrm{VS}}$
We state above 3 core properties of $q\overline{\mathrm{VS}}$ that makes it suitable for measuring cultural diversity in T2I models: 1) prioritizes collections with higher-quality items when other factors are equal, 2) penalizes the duplication of elements, and 3) exhibits flexibility in capturing various aspects of diversity through the selection of an appropriate similarity kernel. The proof of the quality-awareness property is immediate following the definition of $q\overline{\mathrm{VS}}$. See Appendix K for a proof of duplication scaling.
@@ 5.2 Experimental Setup
We discuss the experimental pipeline: 1) Prompting and Seeding: We calculate $q\overline{\mathrm{VS}}$ for 8 images per prompt, matching the typical number of output images of image-generation APIs. To account for variances in prompt wording as well as seed selections, we report scores averaged over 50 repetitions. 2) Mapping Generated Images to cultural artifacts: We map each image to to its most closely resembling artifact from the concept space of the domain. [^3] Not all text-to-image generated images perfectly represent real-world cultural entities. Note that since the prompts focus on global concepts, we obtain the continent, country, and artifact name annotation for each generated image. 3) Computing Vendi Scores: With each generated image linked to its closest artifact, we compute the cultural diversity of the generated outputs using the metric defined in Section 5.1. We expand on the details of each of these steps in Appendix I. Different kernels can capture different aspects of geo-cultural diversity.
@@
  • Continent-level diversity: $w_{1}=1$, $w_{2}=0$, $w_{3}=0$. Considers continent-level similarity.
  • Country-level diversity: $w_{1}=0$, $w_{2}=1$, $w_{3}=0$. Considers country-level similarity.
  • Artifact-level diversity: $w_{1}=0$, $w_{2}=0$, $w_{3}=1$. Only considers distinct artifacts.
@@
  • Uniformly weighted diversity: $w_{1}=1/3$, $w_{2}=1/3$, $w_{3}=1/3$.
Models. We evaluate 4 models across closed-source and open-source model types: 1) Imagen 2, 2) Stable-Diffusion-XL, 3) Playground, and 4) Realistic Vision. More details about the model usage and hyperparameters are provided in Appendix I.2.4.
@@ 5.3 Results
Results in Figure 8 reveals that when prompted with under-specified prompts mentioning general concepts (Fig 2), current T2I models tend to generate artifacts that lack comprehensive geographical representation. This finding aligns with previous observations (Basu et al., 2023), suggesting a bias towards well-represented and popular countries.
@@
Table 4 presents the results for both the average quality score (q) and the diversity component ($q\overline{\mathrm{VS}}$) across different kernels. Playground and Imagen generally achieve the highest quality scores based on the HPS-v2 metric. As anticipated, models exhibit the lowest diversity for $w_{1}=1$, $w_{2}=0$, $w_{3}=0$, which considers only continent-level similarity, due to the limited number of continents. Conversely, $w_{1}=1$, $w_{2}=0$, $w_{3}=0$, focusing solely on artifact diversity, yields the highest scores, reflecting the wide array of potential cultural artifacts. In terms of overall performance, Imagen 2 consistently demonstrates the best $q\overline{\mathrm{VS}}$ scores across different kernels for the Cuisine and Art concepts, whereas SDXL obtains the highest scores for Landmarks concept. Table 5 shows the cultural diversity ($q\overline{\mathrm{VS}}$). Note that the scores across the board are still low, remaining far from the maximum score of 1. Current T2I models fall short of representing the true breadth and richness of global cultural diversity.
Table 5: CD scores across models and concepts using various similarity kernels averaged over 50 repetitions. Imagen 2 (IM) performs best for Cuisine and Art, while SDXL performs best for Landmarks. Importantly, even the best scores are low, indicating significant room for improvement in the cultural diversity of T2I outputs.
@@
6 Discussion and Limitations
To the best of our knowledge, CUBE is the first large-scale cultural competence benchmark for text-to-image models. From our investigations so far, one clear finding stands out: there is yet significant headroom for improvement of global cultural competence in the current generation of text-to-image models — both in terms of awareness and diversity. This seems especially true for the Global South, highlighting the urgency of the need for comprehensive and informative cultural competence testing frameworks. Towards that goal, we have made CUBE dataset and code public [^4] https://github.com/google-deepmind/cube, and encourage its adoption and expansion by the multimodal generative AI community.
Design of this benchmark required many challenging decisions that needed careful thought. The increased coverage of our benchmark as a result of largely automated, scalable approaches - built on curated data sources and existing model capabilities still comes with its own limitations. For instance, existing structured knowledge bases such as WikiData are known to have inherent cultural biases reflecting disparities in global distribution of knowledge production (Callahan & Herring, 2011; Bjork-James, 2021). Therefore, it is important to note that our approach of expanding coverage using existing knowledge bases should be complemented with community based and participatory approaches for richer socio-cultural representation (Alonso Alemany et al., 2023; Dev et al., 2024). Notably, the World Wide Dishes effort (Magomere et al., 2024) builds a dataset of images representing dishes from around the world through a community-led effort that complements our dataset that relies on existing knowledge bases.555www.worldwidedishes.com
Furthermore, even with significant filtering and completion, we expect our data to be noisier than other methods. We have not yet explored the potential utility of our curation method, nor of the dataset itself, to empower other research on evaluating or improving cultural competence in generative models in general. Many aspects of the curation process for CUBE is automated, due to the large scale of the problem — e.g., we rely on image similarity scores and LLM-based selection flows to ground generated images to a specific cultural artifact. This approach risks ingraining biases that may already exist within these tools into the benchmark construction itself. The process of mapping artifacts to country/continent may also introduce biases since the annotator VLM itself may not be aware of several cultural artifacts around the world. On the other hand, human annotation to measure faithfulness and realism is also quite challenging and subjective, as annotators may not be aware of the multitude of representations of their own culture. Even beyond cultural knowledge, different cultures may also have different standards for realism — which could result in mis-calibrated results obtained from human annotations, making it hard to compare across different cultures (e.g., Table 14).
7 Conclusion
We introduced CUBE, a T2I benchmark to assess the cultural competence of T2I models along two crucial dimensions: cultural awareness and cultural diversity. We presented a scalable methodology with a potential to be scaled beyond eigth countries and three concepts considered in this work. Furthermore, we proposed a novel T2I evaluation component: cultural diversity (CD) and measured it using the quality-aware Vendi score. Our comprehensive human evaluation reveals substantial gaps in cultural awareness across cultures and concepts, as well as the gaps in the geo-cultural diversity of model generations. Our correlation analysis reveals a noteworthy trend: while faithfulness and realism exhibit a moderate positive correlation, suggesting they can be improved in tandem, cultural diversity remains weakly correlated to these metrics. By highlighting existing limitations in cultural competence of T2I models, we believe our work contributes to a critical dialogue surrounding the development of truly inclusive generative AI systems.
8 Ethical Considerations
9 Acknowledgements
Appendix A Contributions
Appendix B Correlation Analysis of CD
@@
Appendix C Additional details of CUBE
Below we provide details on the choice of countrues, CUBE-1K dataset breakdown, WikiData root nodes and some technical details for CUBE construction.
C.1 Justification for the choice of countries
We selected eight countries from different geo-cultural regions across continents and the Global South-North divide: Brazil (LatAm), France (Europe), India (SEA), Italy (Europe), Japan (East Asia), Nigeria (SSA), Turkey (Middle East), and USA (North America). Our goal was to choose countries with the largest population in each of these regions, while also taking into account (a) their representation in training data (e.g. Nigeria is “low-resource” whereas USA is “high-resource”), and (b) availability of raters from that region through our vendor. We limited our study to 8 countries because of time and monetary constraints. While we acknowledge that this list of countries is necessarily incomplete, and may result in a biased global sampling, future iterations of this work could include a wider range of countries for a more comprehensive evaluation.
C.2 CUBE-1K statistics
@@ C.3 Wikidata
C.4 Technical Details
We have provided additional technical details on CUBE.
KB Extraction.
We have reported the root nodes for each domain in 8 We iterated for a total of 4 hops beginning from these root nodes as the majority of the artifacts were found in the second and third hop for all the 3 domains we considered. The new artifacts extracted began to plateau after the 4th hop.
Self-Refinement.
We divide the self refinement of the concept space into two steps: 1) Removing noise: An incorrect artifact that does not belong to either that country or cultural concept. We leverage LLMs for this filtering step by asking “Can you classify if the <item> belongs to <country> <concept>? Answer yes or no.”, 2) Adding missing artifacts: We leverage the self-critiquing technique introduced in (Lahoti et al., 2023) by following “critique the response” and “address the critiques and rewrite” steps for each of the artifact lists.
Manual Filtering.
CUBE-1K is intended to serve as a high-quality curated prompt set to represent cultural artifacts selected for relevance and popularity. As noted earlier, we use the local Google Search results as a proxy for popularity within that local context. While this provides us with a broad set of artifacts in each cultural context, some of them may show inflated results because of the commonality of certain words in their names (e.g., “Puri” is the name of a famous temple in India and also a popular dish.). To mitigate this, we used a manual filtering process conducted by annotators from their respective cultures. This process removes any artifacts that may have been artificially boosted by inflated search results. The criteria for manual filtering include visual distinctiveness, alignment with the corresponding geo-cultural category, and the artifact’s popularity within the culture.
Noise in Google Search.
Google Search can potentially inflate the search results due to the presence of common words in the names of cultural artifacts. Here are some examples:
Culture Artifact Search Results Reason
Turkey Van Museum 35.5M Contains "Van"
Japan Japan Monkey Park 75.9M Contains "monkey"
Nigeria Freedom Park 17.8M Contains "freedom"
Table 9: Examples of inflated search results due to the presence of common words in artifact names.
These artifacts are examples of noisy outcomes where the search results were inflated due to the presence of common words.
Annotation Details.
Our annotators were recruited based on the following criteria:
  • Be fluent in English
  • Be diverse in Gender
  • Be from specific countries with familiarity of local culture
@@ C.4 Technical Details
Appendix D Detailed Human Annotation Questions
@@
Table 11: Inter rater reliability for the 3 annotation tasks described in Section D and for all the rater pools across the different locations of our study. For the Cultural Relevance question, we report the observed majority agreement. For both the Faithfulness and the Realism questions, we report the Krippendorff’s $\alpha$.
Model India Japan Italy USA Brazil France Turkey Nigeria
Imagen 98.4 90.8 86.8 95.0 94.3 95.2 68.1 73.6
SDXL 95.2 84.4 82.2 99.2 52.8 96.0 66.4 86.1
@@
Appendix E Inter-Annotator Agreement
We obtain high inter-rater agreement for the question on cultural relevance across all countries (all above 95%; see Table 11), suggesting that the question of whether an image is relevant to a particular culture is a relatively objective task. However, the question of faithfulness and realism yielded moderate to low agreement (especially for France and Turkey) among annotators (measured as Krippendorff’s $\alpha$, which is better suited for Likert scale ordinal values) in line with the relatively more complex and subjective nature of the task (see Table 14 for examples of edge cases).
Appendix F On Realism
Our focus on realism in our evaluation stems from the fact that generative language models are being deployed in products that increasingly shape the discovery of socio-cultural knowledge such as search, online education, and travel planning. In such contexts, cultural awareness is especially important, and realism of generated images is a crucial aspect in this regard. We agree that there may be usage contexts of the T2I models where realism of generated images may not be relevant — for instance, in creative contexts where people use these models to generate photo-realistic images which may be non-realistic in practice (e.g., “photo of Taj Mahal in a desert”). Such generations are not inherently bad, but in contexts where cultural awareness is relevant, our methodology enables the study of cultural awareness of any given model.
Appendix G Background on T2I Evaluation
T2I Evaluation Metrics:
T2I Benchmarks:
Appendix H Background on seed variation in T2I models